# AED: Adaptable Error Detection for Few-shot Imitation Policy

### v1

@@ 5 Pattern Observer (PrObe)
Apart from formulating the AED task, we also develop a tailored method to address it. Our motivation is to detect behavior errors in the space of policy features to acquire additional task-related knowledge. Thus, an adaptable error detector, Pattern Observer (PrObe), is proposed to discover the unexpressed patterns in the policy features extracted from frames of successful or failed states. Even if the visual inputs vary during inference, PrObe leverages the pattern flow and a consistency comparison to alleviate the challenges shown in Figure 1. As a result, it can address the AED task effectively. Furthermore, a rollout augmentation is designed to increase rollout diversity and enhance the method’s robustness, as we now present in detail.
@@ 5.1 Rollout Augmentation
To obtain rollouts with balanced outcomes (successful or failed) and precise frame-level labels, we collect $X^{b}_{succ}$ and $X^{b}_{fail}$ using the agent’s optimal policy $\pi^{\ast}$ with intervention rather than the trained FSI policy $\pi_{\theta}$. However, even if the policy $\pi_{\theta}$ is trained on $X^{b}_{succ}$ (collected from the agent’s optimal policy $\pi^{\ast}$), the trained $\pi_{\theta}$ will still deviate from agent’s $\pi^{\ast}$ due to the limited rollout diversity. In other words, a behavior bias exists between the two policies, which should be reduced. Otherwise, an adaptable error detector trained on original collected $X^{b}_{succ}$ and $X^{b}_{fail}$ will yield a high false-positive during AED inference.
Accordingly, we augment agent rollouts so as to increase rollout diversity and dilute the behavior bias between trained $\pi_{\theta}$ and agent’s $\pi^{\ast}$. Specifically, we iterate each frame (and its label) from sampled agent rollouts and randomly apply the following operations: keep, drop, swap, and copy, with probability 0.3, 0.3, 0.2, and 0.2, respectively. As a result, we inject distinct behaviors into collected rollouts, such as speeding up (interval frames dropped), slowing down (repetitive frames), and not smooth movements (swapped frames). We carefully verify its benefits in Section 6.
@@ 5.2 PrObe Architecture
As depicted in Figure 3, PrObe comprises three major components: a pattern extractor, a pattern flow generator, and an embedding fusion. Since PrObe detects errors through policy features, the history features $f_{h}$ and task-embeddings $f_{\zeta}$ from the trained policy $\pi_{\theta}$ are passed into PrObe.
@@
At first, the pattern extractor aims to gather discriminative information from each embedding of history features $f_{h}$. In more detail, $f_{h}$ is transformed into unit embeddings divided by its L2 norm, mitigating biases caused by changes in visual inputs. Then, a learnable gate composed of a linear layer with a sigmoid function determines the importance of each embedding cell. A Hadamard product between the sign of unit embeddings and the importance scores, followed by an instance normalization (IN), is applied to obtain the pattern features $f_{p}$ for each timestep. PrObe then feeds $f_{p}$ into a flow generator (a LSTM) to generate the pattern flow. Intuitively, the changes in the pattern flow of successful and failed rollouts will differ. On the other hand, the task-embeddings $f_{\zeta}$ are transformed by an IN layer, a linear layer (FC), and a tanh function. This step maps the task-embeddings into a space similar to the pattern flow. Then, a fusion process concatenates the pattern flow and transformed task-embeddings to compute the error predictions $\hat{y}$. This process compares the consistency between the agent rollout and demonstrations and uses this as a basis for determining whether behavior errors occur.
Objectives
PrObe is optimized by one supervised and two unsupervised objectives. First, since accurate frame-level error labels $y$ are available during training, a binary cross-entropy (BCE) objective $L_{cls}$ is used to optimize the error prediction $\hat{y}$. Second, $L_{pat}$ is an L1 loss applied to the pattern features $f_{p}$, which encourages the pattern extractor to generate sparse pattern features, making it easier to observe changes. In addition, a contrastive objective $L_{tem}$ is added to logit embeddings (the embedding before the final outputs) to reinforce the difference between the normal and error states in failed rollouts $X^{b}_{fail}$, i.e., we sample anchors from $X^{b}_{fail}$, and the positive and negative samples will be logits of the same or opposed states from the same rollout.
However, the logits of adjacent frames will contain similar signals since they contain temporal information from the pattern flow, even if behavioral errors have occurred (cf. Appendix’s Figure 7). Blindly forcing the logits of normal and error states to be far away from each other may mislead the model and have a negative impact. Hence, $L_{tem}$ is a novel temporal-aware triplet loss (Schroff et al., 2015) calculated by:
[…]
@@
where $N$ is the number of sampled pairs, and the temporal-aware margin $m_{t}=m*(1.0-\alpha*(d_{ap}-d_{an}))$ will be enlarged or reduced considering the temporal distance of anchor, positive and negative samples. Among them, $m$ is the margin in the original triplet loss, and $d_{ap}$, $d_{an}$ are the clipped temporal distances between the anchor and positive sample and the anchor and negative sample, respectively. Ultimately, the total loss $L_{total}$ combines $L_{pat}$, $L_{tem}$, $L_{cls}$ objectives weighted by $\lambda_{pat}$, $\lambda_{tem}$, and $\lambda_{cls}$.
@@ 6 Experiments
Our experiments aim to verify the following questions: (1) Is learning how to discern patterns in the space of policy features actually beneficial for addressing the AED task? (2) Is our proposed PrObe highly general, robust and interpretable? (3) Is the proposed AED task practical and helpful to the research community?
@@ 6.1 Experimental Settings
Existing robotic manipulation benchmarks (James et al., 2020; Fu et al., 2020; Li et al., 2022; Gu et al., 2023) do not suit our needs due to the novel challenges posed by the AED task. We design seven multi-stage FSI tasks (six indoor and one factory scene) that contain the challenges introduced in Section 3. Detailed task descriptions, schematic, the number of environments, and distractors are shown in the Appendix’s Section D. We build the tasks on Coppeliasim software (Robotics, ) and use Pyrep (James et al., 2019) toolkit as a coding interface.
@@
To verify if the AED methods can handle various policies’ behaviors, we select three representative DC policies to perform FSI tasks, which are implemented following the descriptions in Section 3 and use the same feature extractor and actor architecture. The main difference between them is how they extract task-embeddings from expert demonstrations. NaiveDC (James et al., 2018) [^2] TaskEmb $\rightarrow$ NaiveDC, avoiding confusion w/ AED baselines. concatenates the first and last frames of demonstrations and averages their embeddings as task-embeddings; DCT (Dance et al., 2021) [^3] DCRL $\rightarrow$ DCT (transformer), since we don’t use RL training. performs cross-demonstration attention and fuses them at the time dimension; SCAN (Yeh et al., 2022) computes the tasks-embeddings using stage-conscious attention to locate the critical frames in demonstrations.
@@
PrObe is compared with four error detection methods that possess different characteristics. Unlike PrObe, all methods are policy-independent (i.e., containing their own encoders) and use the same encoder architecture. We provide the detailed baseline introduction in Section C of Appendix and briefly describe them below: (1) SVDDED, an one-class SVM modified from Ruff et al. (2018), is a single-frame DC method trained with only successful rollouts. (2) TaskEmbED, a single-frame DC method trained with successful and failed rollouts, uses the NaiveDC (James et al., 2018) policy’s strategy to extract task-embeddings when predicting errors. (3) LSTMED is a LSTM model (Hochreiter & Schmidhuber, 1997) that predicts errors without demonstration information. (4) DCTED leverages the DCT (Dance et al., 2021) policy’s strategy to predict errors.
Figure 4: AED methods’ performance on our seven challenging FSI tasks. The value under the policy indicates the policy’s success rate in that task. The ranges of the two metrics [$\uparrow$] are both 0 to 1. Our PrObe gets the best 13 AUROC and 15 AUPRC scores out of 21 cases (top of each column), demonstrating its superiority and robustness. We refer to the Appendix’s Table 8 for detailed values.
@@
The area under the receiver operating characteristic (AUROC) and the area under the precision-recall curve (AUPRC), two conventional threshold-independent metrics in error (or anomaly) detection literature, are reported. We linearly select 5000 thresholds spaced from 0 to 1 (or SVDDED’s outputs) to compute the two scores. The accuracy function (cf. Appendix’s Section E.1) determines the error prediction results (i.e., TP, TN, FP, FN) at the sequence-level so that we further conduct and illustrate the timing accuracy experiment in Figure 5.
@@ 6.2 Analysis of Experimental Results
Main experiment
Our main experiment aims to answer the first two questions. We let three policies perform all seven FSI tasks and record their rollouts’ observations and outcomes (success or failure). Then, AED methods infer the error predictions and report values on the two metrics (third stage in Figure 2). It is worth noting that policies behave differently, making it difficult for AED methods to effectively detect behavioral errors in all policies. Furthermore, in our experiments, AUPRC has more reference value because the results of policy rollout are often biased (e.g., many successes or failures), AUPRC is less affected by the class imbalance of outcomes.
@@
From Figure 4, PrObe effectively reports the inconsistency between rollout frames and demonstrations (i.e., behavior errors) in novel environments, even identifying some errors not included in the training data. Specifically, it achieves the best 13 AUROC and 15 AUPRC results out of 21 cases, showcasing its superiority among AED methods and its robustness across different policies. Besides, we investigated the cases where PrObe exhibited suboptimal performance and found that they can be grouped into two types: (1) Errors all occur at the same timing (e.g., at the end of the rollout) and can be identified without reference to the demonstrations (e.g., DCT policy in Organize Table task). (2) Errors are not as obvious compared to the demonstrations, and there is still a chance to complete the task, such as the drawer being closed to a small gap (SCAN and DCT policies in Close Drawer task). In these cases, the pattern flow will not change enough to be recognized, resulting in suboptimal results for PrObe.
@@
Our main experiment examines whether AED methods can accurately report behavior errors when receiving complete policy rollouts. To further verify if they identify errors at the right timing, we manually annotated a subset of rollouts (SCAN policy in Pick &\&& Place task) and visualized the raw outputs of all AED methods. SVDDED outputs the embedding distance between observation and task-embedding, while others compute the probabilities, as shown in Figure 5. In the successful rollout case, both our PrObe and LSTMED stably recognize the states as normal, while other methods misidentify and increase the probability that inputs may be the error states. In the failed rollout case, PrObe detects the error after it has persisted for a while ($<$ 5 frames) and significantly increases the probability. We attribute this short delay to the fact that the pattern flow takes a short period to cause enough change. Even so, PrObe is the one that detects the error at the closest timing; other methods raise the probability either too early or too late. We emphasize that this situation is ubiquitous and not a deliberate choice.
@@
To analyze whether the learned embeddings possess discernible patterns, we initially extract the same 20 rollouts from the annotated dataset above using three AED methods to obtain the embeddings. Subsequently, we present the t-SNE transform (van der Maaten & Hinton, 2008) on these embeddings in Figure 6. Obviously, the embeddings learned by TaskEmbED and LSTMED are scattered and lack an explainable structure. In contrast, our PrObe’s embeddings exhibit characteristics of task progress and homogeneous state clustering, i.e., the states with similar properties (e.g., the beginnings of rollouts, the same type of failures) are closer. This experiment supports the hypothesis that PrObe can learn implicit patterns from the policy’s feature embeddings.
@@
To further verify the second question, we conduct and summarize comprehensive ablations in the Appendix. First, Figure 11 indicates that the proposed rollout augmentation (RA) strategy can increase the rollout diversity and benefits AED methods with temporal information. Second, Figure 12 proves that PrObe’s design can effectively improve performance. Third, Figure 13 shows PrObe’s performance stability by executing multiple experiment runs on a subset of tasks and computing the performance variance. Fourth, we examine how AED methods’ performance is influenced when receiving demonstrations with distinct qualities, as presented in Table 6. Regarding the third question, we provide the pilot study’s result on error correction in Table 7 to demonstrate the AED task’s practicality. Please refer to the corresponding paragraphs in the Appendix for exhaustive versions.
@@ 7 Conclusion
Impact Statement
Our work aims to address a critical learning task (i.e., error detection of policy behaviors), accelerating the development of safety-aware robotic applications in real-world scenarios. We feel that our research will not have any potential negative impacts or ethical concerns for society.
@@
We present more details in this appendix, following a similar flow to the main paper. First, Section A introduces the few-shot imitation (FSI) policies. Then, more our PrObe’s details are provided in Section B. Besides, we introduce the implemented AED baselines in Section C. Next, the designed multi-stage FSI tasks are exhaustively described in Section D, including task configurations, design motivations, and possible behavior errors. In addition, more experimental results are shown in Section E to support our claims. At last, we responsibly present the limitations of this work in Section F.

### v2

@@ 5 Pattern Observer (PrObe)
To address the AED task, we develop a rollout augmentation approach and a tailored AED method. The rollout augmentation aims to increase the diversity of collected rollouts and prevent AED methods from being overly sensitive to subtle differences in rollouts. Regarding the AED method, our insight is to detect behavior errors from policy features that contain task-related knowledge, rather than independently training an encoder to judge from visual observations alone. Thus, we propose Pattern Observer (PrObe), which discovers the unexpressed patterns in the policy features extracted from frames of successful or failed states. Even if the visual inputs vary during inference, PrObe leverages the pattern flow and a consistency comparison to effectively alleviate the challenges posed by the AED task.
@@ 5.1 Rollout Augmentation
To ensure a balanced number of successful and failed rollouts (i.e., $X^{b}_{succ}$ and $X^{b}_{fail}$), along with precise frame-level labels, we gather them using the agent’s optimal policy $\pi^{\ast}$ (with intervention). However, even if the policy $\pi_{\theta}$ is trained on $X^{b}_{succ}$, it inevitably diverges from agent’s optimal policy $\pi^{\ast}$ due to limited rollout diversity. Therefore, AED methods trained solely on the raw collected agent rollouts will be too sensitive to any subtle differences in the trained policy’s rollouts, leading to high false positive rates.
Accordingly, we augment agent rollouts so as to increase their diversity and dilute the behavior discrepancy between trained $\pi_{\theta}$ and agent’s $\pi^{\ast}$. Specifically, we iterate through each frame and its label from sampled agent rollouts and randomly apply the following operations: keep, drop, swap, and copy, with probabilities of 0.3, 0.3, 0.2, and 0.2, respectively. This process injects distinct behaviors into the collected rollouts, such as speeding up (interval frames dropped), slowing down (repetitive frames), and non-smooth movements (swapped frames). We demonstrate that this approach can contribute to AED methods, as shown in Figure 11 in the Appendix.
@@ 5.2 PrObe Architecture
As depicted in Figure 3, PrObe comprises three major components: a pattern extractor, a pattern flow generator, and an embedding fusion. First, the pattern extractor take as input the history features $f_{h}$ from the trained policy $\pi_{\theta}$, aiming to retrieve discriminative information from each embedding of $f_{h}$. Precisely, the pattern extractor transforms $f_{h}$ into unit embeddings through division by its L2 norm, thus mitigating biases caused by changes in visual inputs (first challenge). Then, a learnable gate composed of a linear layer with a sigmoid function determines the importance of each embedding cell. A Hadamard product between the sign of unit embeddings and the importance scores, followed by instance normalization (IN), is applied to obtain the pattern features $f_{p}$ for each timestep.
Second, PrObe feeds $f_{p}$ into an LSTM (flow generator, third challenge) to generate the pattern flow. Intuitively, the changes in the pattern flow of successful and failed rollouts will differ. On the other hand, the task-embeddings $f_{\zeta}$ extracted from $\pi_{\theta}$ are transformed by an IN layer, a linear layer (FC), and a tanh function, mapping the task-embeddings into a space similar to the pattern flow.
Third, a fusion process concatenates the pattern flow and transformed task-embeddings to compute the error predictions $\hat{y}$. This process is expected to compare the consistency between the agent rollout and demonstrations and uses this as a basis for determining whether behavior errors occur.
Objective Functions
PrObe is optimized by one supervised and two unsupervised objectives. Firstly, a binary cross-entropy objective $L_{cls}$ is employed to optimize the error prediction $\hat{y}$, as frame-level labels $y$ are available during training. Secondly, the L1 loss $L_{pat}$ is applied to the pattern features $f_{p}$, encouraging the pattern extractor to generate sparse pattern features, facilitating the observation of pattern changes (second challenge). Thirdly, a contrastive objective $L_{tem}$, a temporal-aware variant of triplet loss [40], is developed and applied to the logit embeddings to emphasize the difference between the normal and error states in failed rollouts $X^{b}_{fail}$ (third challenge). The idea behind $L_{tem}$ is that adjacent frames’ logits contain similar signals due to the temporal information from the pattern flow, even in the presence of errors (cf. Figure 7 in the Appendix). Blindly pushing the logits of normal and error states far apart could mislead the model and have adverse effects. Therefore, $L_{tem}$ considers the temporal relationship between samples and is calculated as follows:
$L_{tem}=\frac{1}{N}\sum_{i=0}^{N}\max(\|\mathrm{logit}_{i}^{a}-\mathrm{logit}_{i}^{p}\|_{2}-\|\mathrm{logit}_{i}^{a}-\mathrm{logit}_{i}^{n}\|_{2}+m_{t},0).$ (2)
@@
Here, $N$ is the number of sampled pairs, and the temporal-aware margin $m_{t}=m*(1.0-\alpha*(d_{ap}-d_{an}))$ adjusts based on the temporal distance of anchor, positive and negative samples. $m$ represents the margin in the original triplet loss, while $d_{ap}$, $d_{an}$ are the clipped temporal distances between the anchor and positive sample, and the anchor and negative sample, respectively. Ultimately, PrObe is optimized through a weighted combination of these three objectives.
@@ 6 Experiments
Our experiments seek to address the following questions: (1) Is it better to solve the AED task by analyzing discernible patterns in policy features rather than using independent encoders to extract features from visual observations? (2) How do our architecture designs contribute, and how do they perform in various practical situations? (3) Can our AED task be effectively integrated with error correction tasks to provide greater overall contribution?
@@ 6.1 Experimental Settings
Existing manipulation benchmarks [41, 42, 43, 44] cannot be used to evaluate the AED task because they do not effectively represent all challenges posed by the AED task, such as cross-environment training and deployment. Therefore, we have designed seven cross-domain and multi-stage robot manipulation tasks that encompass both the challenges encountered in FSI [11] and our AED task. Detailed task information is provided in Section D of the Appendix, including mission descriptions, schematics, configurations, and possible behavioral errors. Our tasks are developed using Coppeliasim [13], with Pyrep [14] serving as the coding interface.
@@
To assess the ability of AED methods to handle various policy behaviors, we implement three demonstration-conditioned (DC) policies to perform FSI tasks, following the descriptions in Section 3 and utilizing the same feature extractor and actor architecture. The primary difference among these policies lies in how they extract task embeddings from expert demonstrations. NaiveDC [8] [^2] TaskEmb $\rightarrow$ NaiveDC, avoiding confusion w/ AED baselines. concatenates the first and last frames of demonstrations and averages their embeddings to obtain task-embeddings; DCT [1] [^3] DCRL $\rightarrow$ DCT (transformer), since we don’t use RL training. employs cross-demonstration attention and fuses them at the time dimension; SCAN [11] computes tasks-embeddings using stage-conscious attention to identify critical frames in demonstrations. More details can be found in Section A of the Appendix.
@@
In our newly proposed AED task, we compare PrObe with four strong baselines, each possessing different characteristics, as detailed in Section C of the Appendix. Unlike PrObe, all baselines incorporate their own encoder to distinguish errors. We describe their strategies for detecting errors here: (1) SVDDED: A deep one-class SVM [16] determines whether the current observation deviates from the centroid of demonstration frames (single frame, OOD). (2) TaskEmbED: A model [8] distinguishes the consistency between the current observation and demonstrations (single frame, metric-learning). (3) LSTMED: A deep LSTM [45] predicts errors solely based on current rollout history (temporal). (4) DCTED: A deep transformer [1] distinguishes the consistency between the current rollout history and demonstrations (temporal, metric-learning).
Figure 4: Performance comparison of AED methods on seven challenging FSI tasks. The values under each policy indicate its success rate for each task. AUROC[$\uparrow$] and AUPRC[$\uparrow$] scores are listed in the upper and lower rows for each policy, respectively, ranging from 0 to 1. According to the statistics table, PrObe achieves the highest Top 1 counts (15 and 17 out of 21 cases), average ranking, and average performance difference in both metrics, demonstrating its superiority and robustness.
@@
We report the area under the receiver operating characteristic (AUROC) and the area under the precision-recall curve (AUPRC), two conventional threshold-independent metrics in error/anomaly detection literature [46, 47]. To compute these scores, we linearly select 5000 thresholds spaced from 0 to 1 (or SVDDED’s outputs) and apply a sequence-level accuracy function (cf. Section E.1 in the Appendix). Furthermore, we conduct an evaluation to verify if AED methods can identify behavior errors timely, as depicted in Figure 5.
@@ 6.2 Analysis of Experimental Results
Main experiment - detecting policies’ behavior errors
Our main experiment aims to verify whether discerning policies’ behavior errors from their features is an effective strategy. We follow the protocol in Figure 2 to conduct the experiment. Specifically, the FSI policies perform all seven tasks, and their rollouts are collected. Then, the AED methods infer the error predictions for these rollouts. Lastly, we report two metrics based on their inference results. Notably, the results of policy rollout are often biased (e.g., many successes or failures). Thus, AUPRC provides more meaningful insights in this experiment because it is less affected by the class imbalance of outcomes.
@@
Due to the diverse nature of policy behaviors and the varying characteristics of baselines, achieving consistently superior results across all tasks is highly challenging. Nonetheless, according to Figure 4, our PrObe achieved the highest Top 1 counts, average ranking, and average performance difference in both metrics, indicating PrObe’s superiority among AED methods and its robustness across different policies. We attribute this to PrObe’s design, which effectively addresses the challenges in the AED task. The average performance difference is the average performance gain compared to the second best method across cases, calculated as $(\mathrm{score}-\mathrm{second\_best\_score})/\mathrm{second\_best\_score}$.
Additionally, we investigated cases where PrObe exhibited suboptimal performance, which can be grouped into two types: (1) Errors occurring at specific timings (e.g., at the end of the rollout) and identifiable without reference to the demonstrations (e.g., DCT policy in Organize Table). (2) Errors that are less obvious compared to the demonstrations, such as when the drawer is closed with a small gap (SCAN and DCT policies in Close Drawer). In these cases, the pattern flow does not change enough to be recognized, resulting in suboptimal results for PrObe.
@@
Our main experiment examines whether AED methods can accurately identify behavioral errors when provided with complete policy rollouts. To further validate the timing of their predictions, we annotated a subset of rollouts (specifically the SCAN policy in Pick &\&& Place) and visualized the raw outputs of all AED methods. SVDDED outputs the embedding distance between observation and task-embedding, while others compute probabilities, as depicted in Figure 5.
In the successful rollout case, both our PrObe and LSTMED consistently identify states as normal, while other methods misidentify them, increasing the probability that inputs represent error states. Conversely, in the failed rollout case, PrObe detects errors after a brief delay ($<$ 5 frames), significantly elevating the probability. We attribute this short delay to the time it takes for pattern flow to induce sufficient change. Nonetheless, PrObe detects errors with the closest timing; other methods either raise probabilities too early or too late. We emphasize that this phenomenon is common and not a deliberate choice.
@@
To analyze whether the learned embeddings possess discernible patterns, we initially extract the same 20 rollouts from the annotated dataset above using three AED methods to obtain the embeddings. Subsequently, we present the t-SNE transform [48] on these embeddings in Figure 6. Obviously, the embeddings learned by TaskEmbED and LSTMED are scattered and lack an explainable structure. In contrast, our PrObe’s embeddings exhibit characteristics of task progress and homogeneous state clustering, i.e., the states with similar properties (e.g., the beginnings of rollouts, the same type of failures) are closer. This experiment supports the hypothesis that PrObe can learn implicit patterns from the policy’s feature embeddings.
@@
In response to the second question, we summarize comprehensive ablations in the Appendix. First, Figure 11 indicates that the proposed rollout augmentation (RA) strategy increases the rollout diversity and benefits AED methods with temporal information. Second, Figure 12 demonstrates that PrObe’s design effectively improves performance. Third, Figure 13 illustrates PrObe’s performance stability by executing multiple experiment runs on a subset of tasks and computing the performance variance. Fourth, we examine how AED methods’ performance is influenced when receiving demonstrations with distinct qualities, as presented in Table 6. Lastly, we study the influence of viewpoint changes in Table 7. Regarding the third question, we provide the pilot study’s result on error correction in Table 8 to demonstrate the practicality of our AED task. Please refer to the corresponding paragraphs in the Appendix for exhaustive versions.
@@ 7 Conclusion
This appendix presents more details of our work, following a similar flow to the main paper. First, Section A introduces the few-shot imitation (FSI) policies. Then, our PrObe’s details are provided in Section B. Besides, the strong AED baselines are depicted in Section C. Next, the designed multi-stage FSI tasks are exhaustively described in Section D, including task configurations, design motivations, and possible behavior errors. In addition, more experimental results are shown in Section E to support our claims. At last, we responsibly present our work’s limitations in Section F.